# Deep learning reveals key predictors of thermal conductivity in covalent organic frameworks

### v2

@@
Abstract
The thermal conductivity of covalent organic frameworks (COFs), an emerging class of nanoporous polymeric materials, is crucial for many applications, yet the link between their structure and thermal properties remains poorly understood. Analysis of a dataset containing over 2,400 COFs reveals that conventional features such as density, pore size, void fraction, and surface area do not reliably predict thermal conductivity. To address this, an attention-based machine learning model was trained, accurately predicting thermal conductivities even for structures outside the training set. The attention mechanism was then utilized to investigate the model’s success. The analysis identified dangling molecular branches as a key predictor of thermal conductivity, a discovery supported by feature importance assessments conducted on regression models. These findings indicate that COFs with dangling functional groups exhibit lower thermal transfer capabilities. Molecular dynamics simulations support this observation, revealing significant mismatches in the vibrational density of states due to the presence of dangling branches.
keywords:
Covalent Organic Frameworks, Thermal Conductivity, Deep Learning, Attention
Prakash Thakolkaran1$\dagger$ Yiwen Zheng2$\dagger$ Yaqi Guo1 Aniruddh Vashisth2,* Siddhant Kumar1,*
{affiliations}
1Department of Materials Science and Engineering, Delft University of Technology, 2628 CD Delft, The Netherlands
2Department of Mechanical Engineering, University of Washington, Seattle, WA, USA
[…]
*Email: vashisth@uw.edu, sid.kumar@tudelft.nl; equal contribution
@@ Introduction
Figure 1: Overview of this study. A COF structure consists of molecular substructures called linkers and knots arranged in a periodic pattern described by the topology. Through molecular dynamics, we construct a large dataset of 2,471 COFs (with diverse linkers, knots, and topologies) and their corresponding thermal conductivities. We demonstrate that conventional descriptors, such as density, pore size, void fraction, and surface area, fail to predict the thermal conductivity reliably. We employ a machine learning framework based on attention mechanism and transformer architecture to uncover a novel predictor, thus enhancing our understanding of the structure-property relationship of thermal conductivity of COFs.
@@
Covalent organic frameworks [1, 2] (COFs) are an emerging class of nanoporous polymeric materials. Compared to metal-organic framework (MOFs) [3, 4] and zeolites [5], the crystalline backbone of COFs is composed of organic building blocks – known as knots and linkers – connected by strong covalent bonds, thereby offering higher stability [6]. The crystalline nature of COFs, along with their high surface areas, tunable pore sizes, and functionalizable organic linkers, make them exceptionally suited for a wide range of promising applications. These applications include (but are not limited to) photoconductivity [7, 8], chemo-sensing [9, 10], catalysis [11], drug delivery [12, 13], thermoelectrics [14, 15], semiconductors [16], gas storage and separation [17, 18].
@@
A key property dictating the applications of COFs is thermal transfer. For example, low thermal conductivity is desired for thermoelectrics to maintain large internal thermal gradients and increase efficiency [19]. On the other hand, a high thermal conductivity is desired for gas adsorption and separation where efficient thermal transfer is important for the longevity and stability of the nanoporous membranes [20]. COFs are promising nanoporous candidates for such applications as the ability to design their crystalline topology (e.g., lattice type, pore size) and chemistry via choice of knots and linkers opens up a large and diverse space of thermal conductivities [21, 22, 23, 1, 2]. Therefore, it is of great interest to gain a deeper understanding of the thermal transfer mechanism and to obtain strong structure-property trends, which in turn enables the application-specific design of COFs.
@@
Two approaches can be used to elucidate the thermal structure-property relationships of COFs: experimental trial-and-error methods which involve synthesis and characterization [24, 25], and computational high-throughput screening which relies on first-principle calculations [26, 27, 28]. With COFs offering practically an unlimited design space, an experimental trial-and-error approach to screen new COF candidates (including developing new synthesis routes for each candidate) is prohibitively inefficient. Moreover, experimentally synthesized COFs will always contain crystalline defects [29, 30], which greatly influence the thermal conductivity. This makes it difficult to relate the thermal transfer mechanisms to the geometrical and chemical make-up of a COF structure. On the contrary, molecular dynamics (MD) simulations [31] offer a high-throughput virtual screening alternative to lab-based experiments. However, despite recent advances in computing hardware, even virtual screening can be prohibitively inefficient for a large design space. This is highlighted by the example that generating a dataset of just 2,471 two-dimensional COFs and their thermal conductivity for this study required 1.3 million CPU-hours in cloud computing (translating into approximately four months and \$70,000 in time and cost, respectively). This underscores the need for efficient structure-property maps that are accurate, generalizable, and interpretable.
@@
Recently, Islamov et al. [32] performed a high-throughput screening study of over 10,000 MOF structures and their thermal conductivity ($\kappa$) using MD. They found that the majority of the MOFs possess $\kappa<1\ \text{W}\text{m}^{-1}\text{K}^{-1}$ with a few exceptions possessing ultra-high thermal conductivity ($\kappa>10\ \text{W}\text{m}^{-1}\text{K}^{-1}$). While COFs are generally more thermally stable (owing to the strong covalent bonding of the building blocks), a recent study by Thakur et al. [33] on high-throughput screening study on over 10,000 COFs demonstrated that, generally, COFs also possess a similar range of $\kappa$, with the majority of the structures exhibiting $\kappa<1\ \text{W}\text{m}^{-1}\text{K}^{-1}$. Interestingly, MOFs and COFs possess similar structure-property trends. (i) Increasing the pore size is mostly sufficient to achieve a low thermal conductivity. (ii) Additionally, one can increase the void fraction, the surface area, or include heavy atoms to further decrease thermal conductivity. (iii) However, to increase the thermal conductivity, there are many factors that need to align. Islamov et al. [32] demonstrated that while lowering the pore size and increasing the density results in a higher ceiling of attainable thermal conductivity, other factors, such as the topology, mass-mismatch, and linker length also need to be accounted for. Thakur et al. [33] came to similar conclusions alongside the observation that aligning the polymeric chains in the structure to heat flow direction raises the ceiling of attainable thermal conductivity. However, these hand-crafted guidelines and ad hoc correlations are insufficient for developing an accurate and generalizable predictive model that captures the thermal conductivity structure-property relationships of COFs. Our study shows that no combination of commonly used descriptors consistently predicts thermal conductivity. Consequently, a definitive method for identifying COFs with tailor-made thermal properties for various applications remains elusive.
@@
To address this knowledge gap, we turn to machine learning (ML) with an emphasis on interpretability and explainability. Previous efforts in ML-assisted design for thermally conductive organic materials include the discovery of polymers with high $\kappa$ through a hierarchical feature selection process [34], a reinforcement learning approach using SMILES-based representations [35], and methods leveraging molecular fingerprints as input features, such as fine-tuning a pre-trained regression model followed by screening [36], training a neural network [37], and employing an active learning approach [38]. Hu et al. [39] provide a review of recent efforts and outlook on ML for thermally conductive organic materials. For the ML modeling of structure-property maps of porous polymers (including but not limited to COFs) there are primarily two routes:
@@
  • high interpretability, limited accuracy: using hand-crafted and pre-extracted high-level of the crystalline network (e.g., pore size, density, surface area, atomic composition) as input to classical regression algorithms [40, 41, 42];
@@
  • high accuracy, limited interpretability: using information-rich but raw graph representation of the crystalline network (where nodes denote atoms with features such as position and atom type, and edges denote bonds with the bond order as the feature) as input to end-to-end deep learning-based regression algorithms [43, 44, 45, 46].
@@
Here, we bridge the two routes to develop an accurate predictive model using deep learning while also combining interpretability insights from deep learning with prior knowledge and descriptors to explain the thermal conductivity structure-property relationships of COFs (see Figure 1 for an overview).
@@ Results
We create a labelled dataset of COFs and their thermal conductivities by selecting 2,471 two-dimensional COFs from the unlabelled dataset of Mercado et al. [47] The COF structures in this design space are made up of one type of nodal linker (also referred to as knot) and one type of connecting linker. The sampled subset consists of COFs with 104 different linkers arranged in 25 different topologies (e.g., honeycomb, kagome, square lattice, etc.), and bonded by four different types of covalent linkages (i.e., carbon-carbon, amine, amide, and imine). We conduct non-equilibrium molecular dynamics (NEMD) simulations to calculate thermal conductivities in two orthogonal in-plane directions (see Supporting Information Section 1.1 for more details). In Supporting Information Section 1.2, we show that there is minimal in-plane anisotropy in thermal conductivity across the dataset. Therefore, we utilize average thermal conductivity in both directions, denoted as $\kappa$, as the quantity of interest to explore the structure-property relationships of COFs. In the following and in Figure 2, we present the observations from data analysis of the key descriptors of COFs and thermal conductivities in our labelled dataset. The distribution of all features and their correlations are detailed in Section 1.3 of the Supporting Information.
@@
Figure 2: Distribution of (a) $\kappa$ versus density (with the color indicating data count per bin), (b) $\kappa$ versus largest pore diameter (LPD), (c) $\kappa$ versus void fraction, and (d) $\kappa$ versus gravimetric surface area (GSA). Also shown are four COF structures (i-iv) with contrasting properties, marked by green triangles in the plots above. The $r$-value indicates the Pearson correlation coefficient.
@@
Density: Figure 2a illustrates the distribution of $\kappa$ vs. density. There is a large spread of densities within the dataset and the majority of the COFs possess a density of 0.45-0.55 $\mathrm{g\ cm^{-3}}$. Moreover, in accordance with previous findings [33], we observe an increasing trend in $\kappa$ with increasing density. The Pearson correlation coefficient between $\kappa$ and density is $r=0.594$, suggesting a moderate positive relationship. However, increasing the density increases the ceiling of attainable $\kappa$, but does not guarantee a high value.
@@
Pore size: The pore size of any COF is represented by the largest pore diameter (LPD). Figure 2b shows that LPD in the dataset is spread over a large range between 5 Å and 80 Å and a concentration of structures in the 15-25 Å region. Moreover, there is an inverse relationship between the pore size and the thermal conductivity, as previously discovered by Freitas et al. [48], where the range of achievable $\kappa$ decreases significantly with increasing pore sizes. This is evidenced by the correlation coefficient $r=-0.508$, indicating a moderate negative correlation. Here again, a small pore size does not guarantee a high thermal conductivity. On the other hand, the pore size can be taken as the sole design parameter to effectively reduce the thermal conductivity range (as mentioned in refs. [33, 32]).
@@
Void fraction: Figure 2c shows that the distribution of the void fractions of COFs across the dataset are within the 0.5–0.96 range, with the biggest share of COFs possessing a void fraction of around 0.85 and higher. Furthermore, a high void fraction of around 0.85 delivers both the largest values and range of attainable $\kappa$ within our dataset. This finding is counterintuitive. Void fraction is typically inversely correlated with density (see Supporting Information Section 1.3). Although low density is generally associated with a low $\kappa$ value, as shown in Figure 2a, the unexpected observation is that a high void fraction (which corresponds to low density) surprisingly results in a high $\kappa$ value, as also depicted in Figure 2c. Beyond a favorable range of void fractions, no distinct trend emerges in the relationship with $\kappa$, as the correlation coefficient $r=-0.213$ suggests only a weak negative relationship.
@@
Surface area: Gravimetric surface area (GSA), i.e., surface area per unit mass varies highly across all the COF structures (see Figure 2d) with a larger cluster of points in the 7000-8000 $\mathrm{m^{2}\ g^{-1}}$ range. We also observe that around intermediate GSA of 7,000 $\mathrm{m^{2}\ g^{-1}}$, we have the widest spread and highest values of $\kappa$. The correlation coefficient of $r=-0.147$ indicates a very weak negative relationship, suggesting that GSA is not a significant predictor.
@@
From the above initial analysis, we notice that the geometrical descriptors used here are insufficient to provide solid trends with respect to thermal conductivity. For example, with all the optimal descriptor values, i.e., a large density, a low pore size, a large void fraction, and an intermediate GSA, we are still not guaranteed to obtain a COF with a high $\kappa$. To elucidate this further, we choose four COFs – denoted by (i)-(iv) – and compare them in Figure 2a-d.
@@
COF (i) has the highest thermal conductivity in the dataset with $\kappa=4.025\ \mathrm{Wm^{-1}K^{-1}}$. The structure has a relatively small pore size with 13.71 Å (see Figure 2a), a medium density with 0.928 $\mathrm{g\ cm^{-3}}$ (see Figure 2b), a relatively high void fraction of 0.844 (see Figure 2c), and an intermediate surface area of 7381 $\mathrm{m^{2}g^{-1}}$. In contrast, the second COF (ii) possesses a similar pore size of 9.88 Å, a similar density of 0.938 $\mathrm{g\ cm^{-3}}$, a similar void fraction of 0.869, and a similar surface area of 7171 $\mathrm{m^{2}g^{-1}}$, but exhibits thermal transfer capabilities almost seven times lower than that of (i) with $\kappa=0.603\ \mathrm{Wm^{-1}K^{-1}}$.
@@
The third COF (iii) has a pore size of 33.73 Å in the intermediate-to-high range, a low density at 0.408 $\mathrm{g\ cm^{-3}}$, a high void fraction at 0.932, and a slightly larger surface area of 8741 $\mathrm{m^{2}g^{-1}}$. Following the trends described in literature [33, 48, 49], the geometrical descriptors are not optimal for a high $\kappa$. Nonetheless, COF (iii) has rather high $\kappa=1.96\ \mathrm{Wm^{-1}K^{-1}}$. Lastly, COF (iv) has an exceptionally high pore size at 80.68 Å and a low density of 0.262 $\mathrm{g\ cm^{-3}}$ but high void fraction of 0.958. Due to the extreme geometry, the COF (iv) has a very low thermal conductivity of $\kappa=0.289\ \mathrm{Wm^{-1}K^{-1}}$.
@@
Furthermore, we trained classical ensemble regression models using these descriptors to predict thermal conductivity (detailed results provided in Supporting Information Section 2) and observed that the models yield poor prediction accuracy. This result highlights the limitations of these descriptors in capturing the structure-property relationships underlying thermal conductivity. In summary, the correlation analysis with the previously introduced descriptors and the presented examples illustrate the fact that the thermal conductivity structure-property relationships are complex and call for further examination.
@@
Figure 3: Schematic of PMTransformer model. A sample COF structure is shown on the left. A crystal graph convolutional neural network computes local embeddings of the COF graph, while GRIDAY computes a three-dimensional energy grid of the structure, which becomes the global embeddings. These embeddings are combined and input into the PMTransformer (which includes a transformer encoder with attention mechanism and a prediction head) to predict $\kappa$. On the right, a parity plot compares $\kappa$ predictions with ground truth values. The dashed line represents the ideal line with zero intercept and unit slope.
@@
Our starting point is the deep learning-based Porous Material Transformer (PMTransformer) developed by Park et al. [46] The PMTransformer is a multi-modal transformer [50] model (see Figure 3 for schematic) designed for universal transfer learning for predicting the properties of porous materials, including COFs, MOFs, and zeolites. It receives two distinct sets of inputs: local and global features. Local features, which reflect the chemistry of building blocks and bonds, are derived from a crystal graph convolutional neural network [51] (CGCNN). The CGCNN directly processes the crystal graph, where atoms and bonds are nodes and edges, respectively, and generates embeddings that capture essential chemical details. These embeddings serve as local features in the PMTransformer. Conversely, global features describe crystalline characteristics, including topological and geometric descriptors such as pore size and surface area. These features are derived from three-dimensional energy grids, which are generated by calculating the interaction energy between the material structure and a methane gas molecule at each grid point using the package GRIDAY [52]. Similar to images in Vision Transformers [53], the energy grids are divided into patches and flattened through linear projections. Eventually, both local and global embeddings are input into a transformer encoder, where the attention mechanism [50] helps the model focus on the most relevant parts of the molecular structure for predicting (in our case) the thermal conductivity.
@@
To enable universal transfer learning, the transformer encoder in PMTransformer is pre-trained on an extremely large dataset of 1.9 million hypothetical porous materials to predict easily obtainable yet essential properties such as topology, void fraction, and building block prediction. This approach ensures the encoder captures critical information necessary for accurately predicting other, more complex properties in downstream tasks with much smaller datasets. For more details regarding the pre-training of PMTransfomer, refer to [46].

### v3

@@
Deep learning reveals key predictors of thermal
conductivity in covalent organic frameworks
Prakash Thakolkaran,a Yiwen Zheng,b Yaqi Guo,a Aniruddh Vashisthb,
and Siddhant Kumara
The thermal conductivity of covalent organic frameworks (COFs), an emerging class of nanoporous polymeric materials, is crucial for many applications, yet the link between their structure and thermal properties remains poorly understood. Analysis of a dataset containing over 2,400 COFs reveals that conventional features such as density, pore size, void fraction, and surface area do not reliably predict thermal conductivity. To address this, an attention-based machine learning model was trained, accurately predicting thermal conductivities even for structures outside the training set. The attention mechanism was then utilized to investigate the model’s success. The analysis identified dangling molecular branches as a key predictor of thermal conductivity, leading us to define the dangling mass ratio (DMR), a descriptor that quantifies the fraction of atomic mass in dangling branches relative to the total COF mass. Feature importance assessments on regression models confirm the significance of DMR in predicting thermal conductivity. These findings indicate that COFs with dangling functional groups exhibit lower thermal transfer capabilities. Molecular dynamics simulations support this observation, revealing significant mismatches in the vibrational density of states due to the presence of dangling branches.
footnotetext: a Department of Materials Science and Engineering, Delft University of Technology, 2628 CD Delft, The Netherlandsfootnotetext: b Department of Mechanical Engineering, University of Washington, Seattle, WA, USAfootnotetext: † Supplementary Information available. See DOI: 00.0000/xxxxxxxx.footnotetext: ‡ These authors contributed equally to this work.footnotetext: ¶ These authors contributed equally to this work. E-mail: Vashisth@uw.edu ; E-mail: Sid.Kumar@tudelft.nl
@@ Introduction
Covalent organic frameworks 1, 2 (COFs) are an emerging class of nanoporous polymeric materials. Compared to metal-organic framework (MOFs) 3, 4 and zeolites 5, the crystalline backbone of COFs is composed of organic building blocks – known as knots and linkers – connected by strong covalent bonds, thereby offering higher stability 6. The crystalline nature of COFs, along with their high surface areas, tunable pore sizes, and functionalizable organic linkers, make them exceptionally suited for a wide range of promising applications. These applications include (but are not limited to) photoconductivity 7, 8, chemo-sensing 9, 10, catalysis 11, drug delivery 12, 13, thermoelectrics 14, 15, semiconductors 16, gas storage and separation 17, 18.
@@
A key property dictating the applications of COFs is thermal transfer. For example, low thermal conductivity is desired for thermoelectrics to maintain large internal thermal gradients and increase efficiency 19. On the other hand, a high thermal conductivity is desired for gas adsorption and separation where efficient thermal transfer is important for the longevity and stability of the nanoporous membranes 20. COFs are promising nanoporous candidates for such applications as the ability to design their crystalline topology (e.g., lattice type, pore size) and chemistry via choice of knots and linkers opens up a large and diverse space of thermal conductivities 21, 22, 23, 1, 2. Therefore, it is of great interest to gain a deeper understanding of the thermal transfer mechanism and to obtain strong structure-property trends, which in turn enables the application-specific design of COFs.
@@
Two approaches can be used to elucidate the thermal structure-property relationships of COFs: experimental trial-and-error methods which involve synthesis and characterization 24, 25, and computational high-throughput screening which relies on first-principle calculations 26, 27, 28. With COFs offering practically an unlimited design space, an experimental trial-and-error approach to screen new COF candidates (including developing new synthesis routes for each candidate) is prohibitively inefficient. Moreover, experimentally synthesized COFs will always contain crystalline defects 29, 30, which greatly influence the thermal conductivity. This makes it difficult to relate the thermal transfer mechanisms to the geometrical and chemical make-up of a COF structure. On the contrary, molecular dynamics (MD) simulations 31 offer a high-throughput virtual screening alternative to lab-based experiments. However, despite recent advances in computing hardware, even virtual screening can be prohibitively inefficient for a large design space. This is highlighted by the example that generating a dataset of just 2,471 two-dimensional COFs and their thermal conductivity for this study required 1.3 million CPU-hours in cloud computing (translating into approximately four months and \$70,000 in time and cost, respectively). This underscores the need for efficient structure-property maps that are accurate, generalizable, and interpretable.
Fig. 1: Overview of this study. A COF structure consists of molecular substructures called linkers and knots arranged in a periodic pattern described by the topology. Through molecular dynamics, we construct a large dataset of 2,471 COFs (with diverse linkers, knots, and topologies) and their corresponding thermal conductivities. We demonstrate that conventional descriptors, such as density, pore size, void fraction, and surface area, fail to predict the thermal conductivity reliably. We employ a machine learning framework based on attention mechanism and transformer architecture to uncover a novel predictor, thus enhancing our understanding of the structure-property relationship of thermal conductivity of COFs.
@@
Recently, Islamov et al. 32 performed a high-throughput screening study of over 10,000 MOF structures and their thermal conductivity ($\kappa$) using MD. They found that the majority of the MOFs possess $\kappa<1\ \text{W}\text{m}^{-1}\text{K}^{-1}$ with a few exceptions possessing ultra-high thermal conductivity ($\kappa>10\ \text{W}\text{m}^{-1}\text{K}^{-1}$). While COFs are generally more thermally stable (owing to the strong covalent bonding of the building blocks), a recent study by Thakur et al. 33 on high-throughput screening study on over 10,000 COFs demonstrated that, generally, COFs also possess a similar range of $\kappa$, with the majority of the structures exhibiting $\kappa<1\ \text{W}\text{m}^{-1}\text{K}^{-1}$. Interestingly, MOFs and COFs possess similar structure-property trends. (i) Increasing the pore size is mostly sufficient to achieve a low thermal conductivity. (ii) Additionally, one can increase the void fraction, the surface area, or include heavy atoms to further decrease thermal conductivity. (iii) However, to increase the thermal conductivity, there are many factors that need to align. Islamov et al. 32 demonstrated that while lowering the pore size and increasing the density results in a higher ceiling of attainable thermal conductivity, other factors, such as the topology, mass-mismatch, and linker length also need to be accounted for. Thakur et al. 33 came to similar conclusions alongside the observation that aligning the polymeric chains in the structure to heat flow direction raises the ceiling of attainable thermal conductivity. However, these hand-crafted guidelines and ad hoc correlations are insufficient for developing an accurate and generalizable predictive model that captures the thermal conductivity structure-property relationships of COFs. Our study shows that no combination of commonly used descriptors consistently predicts thermal conductivity. Consequently, a definitive method for identifying COFs with tailor-made thermal properties for various applications remains elusive.
@@
To address this knowledge gap, we turn to machine learning (ML) with an emphasis on interpretability and explainability. Previous efforts in ML-assisted design for thermally conductive organic materials include the discovery of polymers with high $\kappa$ through a hierarchical feature selection process 34, a reinforcement learning approach using SMILES-based representations 35, and methods leveraging molecular fingerprints as input features, such as fine-tuning a pre-trained regression model followed by screening 36, training a neural network 37, and employing an active learning approach 38. Hu et al. 39 provide a review of recent efforts and outlook on ML for thermally conductive organic materials. For the ML modeling of structure-property maps of porous polymers (including but not limited to COFs) there are primarily two routes:
@@
  • high interpretability, limited accuracy: using hand-crafted and pre-extracted high-level of the crystalline network (e.g., pore size, density, surface area, atomic composition) as input to classical regression algorithms 40, 41, 42;
@@
  • high accuracy, limited interpretability: using information-rich but raw graph representation of the crystalline network (where nodes denote atoms with features such as position and atom type, and edges denote bonds with the bond order as the feature) as input to end-to-end deep learning-based regression algorithms 43, 44, 45, 46.
@@
Here, we bridge the two routes to develop an accurate predictive model using deep learning while also combining interpretability insights from deep learning with prior knowledge and descriptors to explain the thermal conductivity structure-property relationships of COFs (see Fig. 1 for an overview).
@@ Results
We create a labelled dataset of COFs and their thermal conductivities by selecting 2,471 two-dimensional COFs from the unlabelled dataset of Mercado et al. 47 The COF structures in this design space are made up of one type of nodal linker (also referred to as knot) and one type of connecting linker. The sampled subset consists of COFs with 104 different linkers arranged in 25 different topologies (e.g., honeycomb, kagome, square lattice, etc.), and bonded by four different types of covalent linkages (i.e., carbon-carbon, amine, amide, and imine). We conduct non-equilibrium molecular dynamics (NEMD) simulations to calculate thermal conductivities in two orthogonal in-plane directions (see Supplementary Information Section 1.1 for more details). The resulting dataset shows a strong skew toward low thermal conductivity values, with the majority of COFs exhibiting $\kappa<1\ \mathrm{Wm^{-1}K^{-1}}$. In Supplementary Information Section 1.2, we show that there is minimal in-plane anisotropy in thermal conductivity across the dataset. Therefore, we utilize average thermal conductivity in both directions, denoted as $\kappa$, as the quantity of interest to explore the structure-property relationships of COFs. In the following and in Fig. 2, we present the observations from data analysis of the key descriptors of COFs and thermal conductivities in our labelled dataset. The distribution of all features and their correlations are detailed in Section 1.3 of the Supplementary Information.
@@
Fig. 2: Distribution of (a) $\kappa$ versus density (with the color indicating data count per bin), (b) $\kappa$ versus largest pore diameter (LPD), (c) $\kappa$ versus void fraction, and (d) $\kappa$ versus gravimetric surface area (GSA). Also shown are four COF structures (i-iv) with contrasting properties, marked by green triangles in the plots above. The $r$-value indicates the Pearson correlation coefficient.
@@
Density: Fig. 2a illustrates the distribution of $\kappa$ vs. density. There is a large spread of densities within the dataset and the majority of the COFs possess a density of 0.45-0.55 $\mathrm{g\ cm^{-3}}$. Moreover, in accordance with previous findings 33, we observe an increasing trend in $\kappa$ with increasing density. The Pearson correlation coefficient between $\kappa$ and density is $r=0.594$, suggesting a moderate positive relationship. However, increasing the density increases the ceiling of attainable $\kappa$, but does not guarantee a high value.
@@
Pore size: The pore size of any COF is represented by the largest pore diameter (LPD). Fig. 2b shows that LPD in the dataset is spread over a large range between 5 Å and 80 Å and a concentration of structures in the 15-25 Å region. Moreover, there is an inverse relationship between the pore size and the thermal conductivity, as previously discovered by Freitas et al. 48, where the range of achievable $\kappa$ decreases significantly with increasing pore sizes. This is evidenced by the correlation coefficient r=0.508r=-0.508, indicating a moderate negative correlation. Here again, a small pore size does not guarantee a high thermal conductivity. On the other hand, the pore size can be taken as the sole design parameter to effectively reduce the thermal conductivity range (as mentioned in refs. 33, 32).
@@
Void fraction: Fig. 2c shows that the distribution of the void fractions of COFs across the dataset are within the 0.5–0.96 range, with the biggest share of COFs possessing a void fraction of around 0.85 and higher. Furthermore, a high void fraction of around 0.85 delivers both the largest values and range of attainable $\kappa$ within our dataset. This finding is counterintuitive. Void fraction is typically inversely correlated with density (see Supplementary Information Section 1.3). Although low density is generally associated with a low $\kappa$ value, as shown in Fig. 2a, the unexpected observation is that a high void fraction (which corresponds to low density) surprisingly results in a high $\kappa$ value, as also depicted in Fig. 2c. Beyond a favorable range of void fractions, no distinct trend emerges in the relationship with $\kappa$, as the correlation coefficient r=0.213r=-0.213 suggests only a weak negative relationship.
@@
Surface area: Gravimetric surface area (GSA), i.e., surface area per unit mass varies highly across all the COF structures (see Fig. 2d) with a larger cluster of points in the 7000-8000 $\mathrm{m^{2}\ g^{-1}}$ range. We also observe that around intermediate GSA of 7,000 $\mathrm{m^{2}\ g^{-1}}$, we have the widest spread and highest values of $\kappa$. The correlation coefficient of r=0.147r=-0.147 indicates a very weak negative relationship, suggesting that GSA is not a significant predictor.
@@
From the above initial analysis, we notice that the geometrical descriptors used here are insufficient to provide solid trends with respect to thermal conductivity. For example, with all the optimal descriptor values, i.e., a large density, a low pore size, a large void fraction, and an intermediate GSA, we are still not guaranteed to obtain a COF with a high $\kappa$. To elucidate this further, we choose four COFs – denoted by (i)-(iv) – and compare them in Fig. 2a-d.
@@
COF (i) has the highest thermal conductivity in the dataset with $\kappa=4.025\ \mathrm{Wm^{-1}K^{-1}}$. The structure has a relatively small pore size with 13.71 Å (see Fig. 2a), a medium density with 0.928 $\mathrm{g\ cm^{-3}}$ (see Fig. 2b), a relatively high void fraction of 0.844 (see Fig. 2c), and an intermediate surface area of 7381 $\mathrm{m^{2}g^{-1}}$. In contrast, the second COF (ii) possesses a similar pore size of 9.88 Å, a similar density of 0.938 $\mathrm{g\ cm^{-3}}$, a similar void fraction of 0.869, and a similar surface area of 7171 $\mathrm{m^{2}g^{-1}}$, but exhibits thermal transfer capabilities almost seven times lower than that of (i) with $\kappa=0.603\ \mathrm{Wm^{-1}K^{-1}}$.
@@
The third COF (iii) has a pore size of 33.73 Å in the intermediate-to-high range, a low density at 0.408 $\mathrm{g\ cm^{-3}}$, a high void fraction at 0.932, and a slightly larger surface area of 8741 $\mathrm{m^{2}g^{-1}}$. Following the trends described in literature 33, 48, the geometrical descriptors are not optimal for a high $\kappa$. Nonetheless, COF (iii) has rather high $\kappa=1.96\ \mathrm{Wm^{-1}K^{-1}}$. Lastly, COF (iv) has an exceptionally high pore size at 80.68 Å and a low density of 0.262 $\mathrm{g\ cm^{-3}}$ but high void fraction of 0.958. Due to the extreme geometry, the COF (iv) has a very low thermal conductivity of $\kappa=0.289\ \mathrm{Wm^{-1}K^{-1}}$.
@@
Furthermore, we trained classical ensemble regression models using these descriptors to predict thermal conductivity (detailed results provided in Supplementary Information Section 2) and observed that the models yield poor prediction accuracy. In summary, the correlation analysis with the previously introduced descriptors and the presented examples illustrate the fact that the thermal conductivity structure-property relationships are complex and call for further examination.
@@
Fig. 3: Schematic of PMTransformer model. A sample COF structure is shown on the left. A crystal graph convolutional neural network computes local embeddings of the COF graph, while GRIDAY computes a three-dimensional energy grid of the structure, which becomes the global embeddings. These embeddings are combined and input into the PMTransformer (which includes a transformer encoder with attention mechanism and a prediction head) to predict $\kappa$. On the right, a parity plot compares $\kappa$ predictions with ground truth values. We report the mean and standard deviation of the goodness-of-fit across five random seeds. The random seeds affect the initialization of the prediction head, while the rest of the model is initialized from the pre-trained PMTransformer weights. The dashed line represents the ideal line with zero intercept and unit slope.
@@
Our starting point is the deep learning-based Porous Material Transformer (PMTransformer) developed by Park et al. 46 The PMTransformer is a multi-modal transformer 49 model (see Fig. 3 for schematic) designed for universal transfer learning for predicting the properties of porous materials, including COFs, MOFs, and zeolites. It receives two distinct sets of inputs: local and global features. Local features, which reflect the chemistry of building blocks and bonds, are derived from a crystal graph convolutional neural network 50 (CGCNN). The CGCNN directly processes the crystal graph, where atoms and bonds are nodes and edges, respectively, and generates embeddings that capture essential chemical details. These embeddings serve as local features in the PMTransformer. Conversely, global features describe crystalline characteristics, including topological and geometric descriptors such as pore size and surface area. These features are derived from three-dimensional energy grids, which are generated by calculating the interaction energy between the material structure and a methane gas molecule at each grid point using the package GRIDAY 51. Similar to images in Vision Transformers 52, the energy grids are divided into patches and flattened through linear projections. Eventually, both local and global embeddings are input into a transformer encoder, where the attention mechanism 49 helps the model focus on the most relevant parts of the molecular structure for predicting (in our case) the thermal conductivity.
@@
To enable universal transfer learning, the transformer encoder in PMTransformer is pre-trained on an extremely large dataset of 1.9 million hypothetical porous materials to predict easily obtainable yet essential properties such as topology, void fraction, and building block prediction. This approach ensures the encoder captures critical information necessary for accurately predicting other, more complex properties in downstream tasks with much smaller datasets. For more details regarding the pre-training of PMTransfomer, refer to 46.